OpenStudy (anonymous):

what is the range of the function y = 6/(X - 12)? thanks.

OpenStudy (anonymous):

Do you know what range is?

OpenStudy (anonymous):

yes, we consider teh values of y.

OpenStudy (anonymous):

will we multiply both sides by x - 12 first?

OpenStudy (anonymous):

In this example, we don't need to.

OpenStudy (anonymous):

And you generally don't need to do so as this is a y=f(x), so keep the left hand side as simple as possible.

OpenStudy (anonymous):

can we do x = 6/y + 12? so, range is (-ing, 0)U(0, inf)?

OpenStudy (anonymous):

which is the same with \[y\in(-\infty,\infty)\]

OpenStudy (anonymous):

you mean, we can include y = 0?

OpenStudy (anonymous):

Yes, when x tends to plus or minus infinity.

OpenStudy (anonymous):

But please be minded that if we are using real analysis, the range will instead read \[y\in (\infty,0)\cup(0,\infty)\]

OpenStudy (anonymous):

i am confused sorry..

OpenStudy (anonymous):

If you use limit, you find that y can be zero if x hypothetically reach infinity. But in reality you can't, so most people will instead exclude zero.

OpenStudy (anonymous):

oh okay, but, if you were asked to write it in interval notation, how will you write it?

OpenStudy (anonymous):

Yours is find enough.

OpenStudy (agent0smith):

As you did earlier, \[ (-\inf, 0) \cup (0, \inf)\] is interval notation.

OpenStudy (anonymous):

oh okay.. thanks..

OpenStudy (anonymous):

it's easier to see it as x=6/y + 12

OpenStudy (anonymous):

as you can see the only value y can't be is 0.

OpenStudy (campbell_st):

the range of this function is 0 to - infinity

OpenStudy (campbell_st):

there is a horizontal asymptote at y = 0

OpenStudy (agent0smith):

@campbell_st you forgot 0 to +infinity...